ANNUAL REPORT OF THE CENTRAL BUREAU FOR ASTRONOMICAL TELEGRAMS

1997

273 Circulars were issued in 1997, six more than in 1991, and making a new record. Just as had been anticipated, comet C/1995 O1 (Hale-Bopp) put on a splendid show for northernhemisphere observers, continuously at first magnitude and brighter from mid-February until early May. With its prominent dust tail generally estimated as attaining a length of 15–20 degrees in the evening sky, it was perhaps seen by more people than any previous comet, and amateur and professional astronomers obtained what must be a record number of color images of any comet. Some 57 of the year’s Circulars contained items on this comet. Among the more intriguing discoveries was the appearance of a narrow tail of neutral sodium in the general direction of the gas tail, the favorable geometry allowing these tails to be well separated from the dust tail when the comet was at its best. In the southern hemisphere, the comet was still being reported as being seen with the naked eye in November.

Contributing more to the record number of Circulars, however, was the appearance of a record number of 140 extragalactic supernovae. Although three of these attained magnitude 14 or brighter, the median magnitude was 21, and several were as faint as magnitude 24.

Observations of γ-ray bursts also “came of age”, with 109 Circulars containing items on as many ten new such events recorded during the year (and continuing coverage of the July 1996 event). In several cases optical counterparts were found, and in some cases detections at radio wavelengths were also reported.

The number of reports of comets during the year was significantly boosted by the recognition of as many as 36 comets in data (in several cases from 1996) obtained with the LASCO white-light coronagraphs aboard the Solar and Heliospheric Observatory (SOHO) spacecraft. None of these objects was observed from the ground, and all but two of them seem to have been members of the Kreutz sungrazing group that did not survive perihelion passage.

Two distant satellites of Uranus were discovered, the first “irregular” satellites known to orbit that planet, but showing now that all four giant planets are accompanied by both irregular and regular satellites and at least segments of rings.

A significant event for the Central Bureau was the decision, timed to coincide with the meeting of IAU Commission 6 in Kyoto on August 23, to make the Circulars freely available in the World Wide Web. The Circulars are posted there after some delay, ranging from 1 to 72 hours, but this is being done both as a genuine service to the international astronomical community and to put an end to the illicit electronic postings of the Circulars that have gone on for several years anyway. Nevertheless, there is bound to be some impact on the Bureau’s finances (which are partly balanced by the Minor Planet Center), and the full impact of this is not yet clear. Although subscribers have the advantage that they can receive the Circulars passively by e-mail and with no delay, one obvious effect of the free WWW availability has been that their number peaked, at around 600 in mid-year, dropping to 584 in December. The number of subscribers to the printed version of the Circulars continued its steady decline, from 411 in January to 370 in December.

Most of the year’s IAU Circulars were prepared by Bureau Associate Director Daniel W. E. Green. The remainder were generally by the undersigned, with a few by summer student Timothy B. Spahr. Responsibility for the administrative work of the Bureau has been shared by Donna Thompson and Muazzez Lohmiller. Gareth Williams has again very effectively maintained the Bureau’s presence on the World Wide Web (URL http://cfawww.harvard.edu/iau/cbat.html). He, Spahr and student Jen Owens also put some effort into computer-scanning and preparing pre-1982 Circulars for inclusion there.

Brian G. Marsden Director of the Bureau